### v2

# Sparse R-CNN OBB: Ship Target Detection in SAR Images Based on Oriented Sparse Learnable Proposals

Abstract
We present Sparse R-CNN OBB, a novel framework for the detection of oriented objects in SAR images leveraging sparse learnable proposals. The Sparse R-CNN OBB has streamlined architecture and ease of training as it utilizes a sparse set of 300 proposals instead of training a proposals generator on hundreds of thousands of anchors. To the best of our knowledge, Sparse R-CNN OBB is the first to adopt the concept of sparse learnable proposals for the detection of oriented objects, as well as for the detection of ships in Synthetic Aperture Radar (SAR) images. The detection head of the baseline model, Sparse R-CNN, is re-designed to enable the model to capture object orientation. We train the model on RSDD-SAR dataset and provide a performance comparison to state-of-the-art models. Experimental results show that Sparse R-CNN OBB achieves outstanding performance, surpassing most models on both inshore and offshore scenarios. The code is available at: www.github.com/ka-mirul/Sparse-R-CNN-OBB.
Index Terms—  Convolutional neural network, deep learning, oriented ship detection, sparse learnable proposals, synthetic aperture radar (SAR).
1 Introduction
Synthetic Aperture Radar (SAR) is a powerful active microwave imaging technology that functions reliably under all weather conditions and at any time of day. Among its various applications in maritime scenarios, ship detection is one of the most prevalent, applicable for both civilian and military purposes.
The ship detection task involves localizing ships within an image and determining their classes. In traditional SAR ship detection systems, the process generally involves three key stages: pre-processing, candidate extraction, and discrimination. The pre-processing step may involve enhancing the quality of the input image, reducing noise, and conducting land-sea segmentation to avoid extracting non-ocean targets. In candidate extraction stage, potential ship targets are identified. Methods based on the Constant False Alarm Rate (CFAR) technique are commonly used for this task. CFAR-based detectors are adaptive threshold systems that estimate sea clutter statistics surrounding a potential target using an assumed background probability density function. This estimation enables them to maintain a consistent, acceptable probability of false alarm (PFA) [1]. Finally, the discrimination stage ensures that only genuine ships are retained for further analysis. Although traditional detectors generally deliver satisfactory performance, their accuracy is prone to variability due to their heavy reliance on hand-crafted parameters assigned at each stage. Furthermore, the lack of end-to-end design in these approaches increases design complexity, making both training and adaptation to other datasets more challenging.
With the advancement of Convolutional Neural Networks (CNNs), there is a growing trend toward their use in developing Oriented Bounding Box (OBB) detectors for oriented ship detection in SAR images. Generally, oriented SAR ship detectors draw inspiration from detectors for common oriented objects which can be categorized into three classes, one-stage, two-stage, and anchor-free detectors. One stage and two-stage detectors are also known as anchor-based approaches due to their dependence on pre-defined anchor boxes. The aim of model training is to regress these boxes to accurately align with object boundaries. Unlike two-stage detectors that employ a small network to first extract the object candidates (also called proposals), one-stage detectors directly regress the anchor boxes. One-stage detectors include R-RetinaNet [2], S2ANet [3], and R3Det [4], while two-stage detectors include Gliding Vertex [5], Oriented RCNN [6], and ReDet [7]. Additionally, CFA [8] and BBAV [9] are anchor-free detectors that perform predictions directly without relying on predefined anchor boxes.
Overall, anchor-based approaches remain favorable due to their superior performance. However, the use of dense anchors introduces persistent challenges such as resulting in redundant outputs and causing model performance to be strongly dependent on the initial anchor configurations. These challenges have led to the utilisation of fewer anchors, introducing the concept of sparse proposals.
In this work, we introduce Sparse R-CNN OBB, a new family of Region-based Convolutional Neural Networks (R-CNN) utilizing rotated sparse proposals, for the detection of oriented ships in SAR images. Sparse R-CNN OBB eliminates the need for hundreds of thousands of anchors. Instead, it utilizes only 300 identically generated learnable proposals. Our main contributions in this article are twofold. First, we developed Sparse R-CNN OBB, first of its kind to adopt sparse learnable proposals for the detection of oriented objects as well as for the detection of ships in SAR images. Second, we fine-tuned the model on oriented SAR ship dataset, RSDD-SAR [10], and provide comparisons to state-of-the-art models.
The remainder of this article is organized as follows, Section 2 explains the concept of sparse learnable proposals. Section 3 provides the implementation details of Sparse R-CNN OBB . Experimental details are provided in Section 4, while the experimental results and performance comparison to other models are covered in Section 5. Finally, Section 6 provides concluding remarks.
2 Related Work
2.1 Sparse Learnable Proposals
The concept of utilizing sparse proposal detectors has been explored in prior research [11, 12, 13, 14], with G-CNN [11] serving as a pioneering approach in this domain. Instead of relying on a proposal generator, G-CNN initiates with a multi-scale grid comprising 180 fixed bounding boxes. These initial boxes are iteratively refined by a regressor that adjusts their position and scale to better align with the objects in the image. At the time of its introduction, G-CNN achieved performance comparable to Fast R-CNN [15], which operates with approximately 2,000 bounding boxes. However, G-CNN lags behind the performance of its successor, Faster R-CNN [16].
DETR [12] advanced the field by proposing a transformer-based architecture that redefines object detection as a set prediction problem. DETR employs a fixed set of 100 learned object queries to simultaneously capture object relationships and global image context, enabling direct parallel predictions. Despite its conceptual elegance, DETR relies on dense interactions between object queries and global image features, resulting in slow convergence during training and hindering the development of a fully sparse detection pipeline
Sparse R-CNN [13] is the first framework to introduce a fully sparse proposal paradigm for object detection. Instead of relying on extensive candidate boxes or dense interactions with global image features, Sparse R-CNN employs a compact set of learnable proposals, known as sparse learnable proposals, comprising both bounding box coordinates and feature embeddings. These proposals interact directly with Region of Interest (RoI) features in a one-to-one manner, eliminating the need for global context processing. This fully sparse interaction significantly reduces computational complexity while achieving efficient and precise object detection. The sparse learnable proposals in Sparse R-CNN provide key advantages over traditional dense proposals. They eliminate the need for a region proposal network (RPN) during training and non-maximum suppression (NMS) during inference as densely-overlapping predictions are inherently avoided. This results in a simplified architecture comprising only a backbone feature extractor, an interaction module, and classification/regression heads
The Sparse Anchoring Network (SAN) proposed in [14] represents another effort to incorporate sparse proposals by generating a reduced set of region proposals. However, SAN operates as a region proposal network (RPN) within the R-CNN framework, preventing it from achieving the simplicity and efficiency of the fully sparse design of Sparse R-CNN.
While the architecture of Sparse R-CNN OBB  will be introduced at a later stage, it is worth noting here that the training regime of Sparse R-CNN OBB is similar to the second stage of Faster R-CNN [16] training. After feature maps generated by the backbone network, the Region of Interest (RoI) features are extracted through RoIPooling operation on one of the map. The key difference is that, unlike Faster R-CNN, which directly feeds RoI features to the classification and regression branches, Sparse R-CNN OBB first interacts them with proposal features through dynamic head.
2.2 Dynamic Head
The dynamic head facilitates the interaction between the proposal features and pooled RoI features, yielding the final object features. This interaction is realized through two consecutive 1x1 convolution layers with ReLU activation functions. The resulting object features are then passed to the regression and classification branches to obtain box offsets and class logits for the proposals. The box offsets are then used to update proposal box parameters.
In the original implementation of Sparse R-CNN, a stack of dynamic head was utilized to enhance the model performance. In this schema, the updated box parameters and the resulted object features are fed into the subsequent head, forming an iterative refinement process. Following this strategy, our proposed Sparse R-CNN OBB employs the same approach by utilizing a stack of six dynamic heads.
3 Sparse R-CNN OBB
This section presents the details of the proposed model along with the strategy used to incorporate the orientation into the model.
3.1 Backbone
Sparse R-CNN OBB employs a ResNet-50 backbone with a Feature Pyramid Network (FPN) for multi-scale feature fusion, referred to as ResNet-50-FPN. As depicted in Fig. 2, residual blocks of ResNet-50 (res2, res3, res4, res5) first output feature maps C2, C3, C4, and C5, which are then passed to the FPN. Afterwards, to standardize their channel dimensions (256, 512, 1024, 2048), a 1×1 lateral convolution is applied to align all these maps to 256 channels. The aligned maps are further iteratively fused in a top-down manner, with smaller maps upsampled to match the larger ones. Finally, a final 3×3 convolution refines the fused maps, producing outputs P2, P3, P4, and P5.
3.2 Rotated Sparse Learnable Proposals
We exploit the concept of sparse learnable proposals, originally built for detecting non-rotated objects, to further accommodate the detection of oriented objects. In our pipeline, each proposal contains 256-dimensional learnable proposal features and 5-dimensional box parameters $x_{p}$, $y_{p}$, $w_{p}$, $h_{p}$, $\theta_{p}$. As depicted in Fig. 1, the ($x_{p}$, $y_{p}$) represents the center point of the proposal while ($w_{p}$, $h_{p}$, $\theta_{p}$) respectively represent the width, height, and orientation of the proposal.
Fig. 1: Oriented bounding box representation used in the proposed Sparse R-CNN OBB.
Fig. 2: Flowchart of the proposed Sparse R-CNN OBB.
To accommodate the embedding of the orientation parameter $\theta_{p}$, we re-designed the baseline model [13] with two significant adjustments. First, in the RoI pooling stage, instead of using standard RoIAlign [17], we used Rotated RoIAlign (R-RoIAlign) [6] on the feature maps outputted by ResNet-50-FPN. This is intended to better capture the feature of the targets enclosed by rotated proposal boxes. Secondly, we modified the structure of the regression layers so that they output five parameters instead of four, representing the offset of each rotated box parameter: $\delta_{x}$, $\delta_{y}$, $\delta_{w}$, $\delta_{h}$, and $\delta_{\theta}$. The updated box parameters are then obtained by following equation:
$\begin{gathered}\hat{x}=x_{p}+\delta_{x}\cdot w_{p}\cdot\cos\theta_{p}+\delta_{y}\cdot h_{p}\cdot\sin\theta_{p},\\ \hat{y}=y_{p}+\delta_{x}\cdot w_{p}\cdot\sin\theta_{p}+\delta_{y}\cdot h_{p}\cdot\cos\theta_{p},\\ \hat{w}=w_{p}\cdot e^{\delta_{w}},\\ \hat{h}=h_{p}\cdot e^{\delta_{h}},\\ \hat{\theta}=\theta_{p}+\delta_{\theta},\end{gathered}$ (1)
where ($\hat{x}$, $\hat{y}$, $\hat{w}$, $\hat{h}$, $\hat{\theta}$) represent the updated proposal box parameters. The schematic of proposed Sparse R-CNN OBB is provided in Fig. 2.
As reported in [13], the initial proposal box parameters ($x_{p}$, $y_{p}$, $w_{p}$, $h_{p}$, $\theta_{p}$) have negligible effect on performance. Therefore, we initialize them by placing boxes at the image center, with width and height set to $\frac{1}{4}$ and $\frac{1}{2}$ of the image size, and an orientation of $-\frac{\pi}{4}$.
3.3 Loss Function
In  Sparse R-CNN OBB training, two types of loss calculations are performed: matching loss and training loss. The matching loss measures the differences between proposals and the ground truth values. The matching cost is formulated as:
$\mathcal{L}=\lambda_{\text{cls}}\cdot\mathcal{L}_{\text{cls}}+\lambda_{\text{L1}}\cdot\mathcal{L}_{\text{L1}}+\lambda_{\text{iou}}\cdot\mathcal{L}_{\text{iou}}$ (2)
In this formulation, the total loss is a weighted sum of the focal loss $\mathcal{L}_{\text{cls}}$, L1 bounding box regression loss $\mathcal{L}_{\text{iou}}$, and IoU loss $\mathcal{L}_{\text{iou}}$, with coefficients $\lambda_{\text{cls}},\lambda_{\text{L1}},\lambda_{\text{iou}}$. The $\lambda_{\text{cls}}$, $\lambda_{\text{L1}}$, and $\lambda_{\text{IoU}}$ coefficients were set to 2.0, 5.0, and 2.0, respectively, following the baseline Sparse R-CNN settings [13]. The training loss, computed only for matched pairs, follows the same structure as the matching loss and is normalized by the number of objects in the batch.
4 Experiment Setup
4.1 Dataset
The Sparse R-CNN OBB is trained and evaluated on the RSDD-SAR dataset [10], specifically designed for tuning oriented SAR ship detector. The dataset contains 7,000 images (512×152 pixels) from TerraSAR-X and Gaofen-3 satellites, with 10,263 annotated ships. It covers spatial resolutions from 2 to 20 meters and includes multiple polarization modes (HH, HV, VH, VV, DH, DV). RSDD-SAR provides both inshore and offshore test sets for evaluating model performance in different environments.
4.2 Evaluation
In this work, average precision ($AP$[18] is utilized as evaluation metric. $AP$ is defined as the area under the PrecisionRecall (PR) curve formulated as:
$AP=\int_{0}^{1}\text{P}(\text{R})\,d(\text{R}).$ (3)
The Precision (P) and Recall (R) are defined as follow:
$Precision=\frac{N_{TP}}{N_{GT}}$ (4)
$Recall=\frac{N_{TP}}{N_{DET}}.$ (5)
Here, $N_{TP}$, $N_{GT}$, and $N_{DET}$ represent the number of true positive predictions, ground truth boxes, and predictions made by the model, respectively. Furthermore, Intersection over Union (IoU) is employed to determine true positive detections. Specifically, we used $AP_{50}$, which requires a minimum IoU score of 0.5 between predicted and ground truth boxes to assign a prediction as a true positive. Additionally, we employed also $AP_{50}Inshore$ and $AP_{50}Offshore$ to evaluate our model in various test sets.
4.3 Hyperparameters and Environment
Sparse R-CNN OBB is set up with weights pre-trained on ImageNet [19]. The model is then trained using the AdamW optimizer [20] with a base learning rate of $7.5\times 10^{-5}$ for 150 epochs, with momentum set to 0.9, weight decay set to $1\times 10^{-4}$. A warm-up strategy is implemented for the first 1000 iterations, with a learning rate 1% of the base rate. The base learning rate is further reduced by a factor of 10 at the 130th and 140th epochs. Training uses a batch size of 8 on two NVIDIA RTX 2080 GPUs with Detectron2 [21] in PyTorch on Ubuntu 22.04, hosted by the Advanced Computing Research Centre’s HPC systems at the University of Bristol.
5 Results and Discussion
In this section, we present the experimental results and compare the performance of Sparse R-CNN OBB  against state-of-the-art (SOTA) models.
5.1 Effect of Number of Proposals
We conducted an experiment by varying the number of proposals to asses their effects on accuracy, model size, training time, and inference speed.
Table 1: Experiment on Number of Proposals
Proposals AP50(%) Model size (M) Training Duration (h) FPS
100 91.02 106.13 11.63 14
200 91.32 106.15 13.34 12
300 91.78 106.18 14.60 11
According to Table 1, increasing the number of proposals marginally enhances performance, but at the cost of longer training times and reduced inference speed. The results also demonstrate that using more proposals does not significantly increase model size. This is because the increase in proposals only adds parameters for storing the 5-dimensional proposal boxes and 256-dimensional proposal features, which are minimal compared to the overall model size.
Considering the above performance, the 300-proposal configuration is set as the default for the model, employed for the rest of the experiments. However, it should be noted that the number of proposals limits the maximum detectable objects and should be adjusted based on the application. For instance, using 100 proposals would be inappropriate for detecting hundreds of objects in a single image.
5.2 Comparison to SOTA Models
In this section, we compare the performance of Sparse R-CNN OBB against state-of-the-art algorithms. As shown in Table 2, Sparse R-CNN OBB outperforms other methods across all metrics, except for $AP_{50}Inshore$. In mixed-scene test sets, reflected by $AP_{50}$, our method demonstrates superior performance, surpassing all state-of-the-art algorithms by a significant margin ranging from 2.51% to 25.16%, with CFA [8] being the closest competitor. Compared to other two-stage detectors, our model demonstrates a superior performance with an improvement of 2.98% over Oriented R-CNN [6], the top-performing algorithm in this category. This performance comparison shows that Sparse R-CNN OBB is robust for detecting ships in typical operational ocean environments, where both inshore and offshore backgrounds may coexist.
Fig. 3: Detection results across different environments, with predictions (magenta) and ground truth (green). The first two left columns show results in inshore regions, while the remaining columns correspond to offshore regions.
Table 2: Performance Comparison on RSDD-SAR Dataset
Type Model AP50 (%) AP50 Inshore (%) AP50 Offshore (%)
Two-stage R-Faster R-CNN 83.29 48.78 90.93
RoI Transformer 88.39 60.83 94.35*
Gliding Vertex 85.55 55.93 91.65
Oriented R-CNN 88.84 65.92 90.21
One-stage R-RetinaNet 66.66 33.20 74.06
S2ANet 87.91 63.27 93.14
R3Det 80.87 56.87 90.16
Anchor-free Polar Encoding 87.31 59.69 90.12
R-FCOS 85.48 50.02 93.09
CFA 89.31* 66.40 90.47
Sparse R-CNN OBB 91.82 66.27* 96.26
Highest value, * Second highest value.
The performance of other models were taken from [10].
In inshore scenes, Sparse R-CNN OBB achieves performance comparable to CFA, with a marginal gap of less than 0.02%. While CFA is tailored for densely packed, arbitrarily oriented objects using convex-hull-based adaptation, Sparse R-CNN OBB is specifically designed for SAR ship detection across both inshore and offshore settings. These results suggest that incorporating inshore-specific refinements could further enhance performance. Nonetheless, Sparse R-CNN OBB remains highly competitive, exhibiting strong robustness and lower false detection rates in complex environments. This result indicates that Sparse R-CNN OBB can better detect ships in complex environments, whereas other methods struggle with higher false detection rates.
Lastly, for offshore scenes, Sparse R-CNN OBB demonstrates exceptional performance, achieving more than 96% accuracy and surpassing other models by at least a 1.91% margin. This results signifies that Sparse R-CNN OBB is more robust against noisy SAR ocean images, which may also be interfered by sea waves and ship-generated wakes.
Finally, Fig. 3 presents sample prediction results from Sparse R-CNN OBB on both datasets across different background scenarios. These results visually confirm that the proposed model accurately detects most ship targets in accordance with their ground truth, demonstrating its robust performance in varying scenarios. We anticipate that our Sparse R-CNN OBB presents a versatile solution for detection of oriented ships, delivering enhanced accuracy while maintaining ease of design and training.
6 Conclusion
In this article, we proposed Sparse R-CNN OBB, a new pipeline for oriented ship detection in SAR imageries based on sparse learnable proposals. The adoption of sparse proposals concept ensures a simple design for our model. Furthermore, the incorporation of rich instance features from learnable proposals improves the model performance.  Sparse R-CNN OBB achieves superior performance, surpassing all state-of-the-art benchmarks on the RSDD-SAR dataset by at least 2.51% in $AP_{50}$.